# From Architectures to Applications: A Review of Neural Quantum States

### v2

@@ 3.3 Dynamics
Regularization schemes for $S$ have been developed, see Sec. 3.1, but these still require a very large number of samples [203] and impact the accuracy [183]. The choice of regularization method impacts the stability of the dynamics [88]. In Ref. [87], this problem is approached by disregarding the contributions to the TDVP of which there is insufficient information available due to constraints of a finite number of samples $N_{MC}$. This still requires the diagonalization of $S$, which has a high computational cost ($O(N_{\bm{\theta}}^{3})$) [203].
@@ 3.5 Open Systems
3.6 Learning from Data
3.6.1 Quantum State Tomography
@@
$\mathcal{M}(\hat{\rho})=\Big{\langle}\hat{U}^{\dagger}\ket{\sigma}\bra{\sigma}\hat{U}\Big{\rangle}_{\sigma}$ (59)
from measurements in the computational basis |$\ket{\sigma}$. The density matrix $\hat{\rho}$ can be approximated by the so-called classical shadow $\mathcal{S}$ consisting of $N_{s}$ samples, with
$\mathcal{S}=\Bigg{\{}\mathcal{M}^{-1}\left(\Big{\langle}\hat{U}_{i}^{\dagger}\ket{\sigma}_{i}\bra{\sigma}_{i}\hat{U}_{i}\Big{\rangle}_{\sigma_{i}}\right)\quad\mathrm{with}\quad i\in 1,\dots,N_{s}\Bigg{\}}.$ (60)
@@
Also in the setting of QST, the choice of NQS depends on the state under consideration, with potential advantages e.g. of autoregressive networks and their perfect sampling [97] or of networks which can represent a high degree of entanglement [136, 254]. Typical architectures are RBMs [241, 61, 255, 37] (see Sec. 2.2), RNNs and transformer networks [136, 254, 123, 37, 256] (see Sec. 2.6), CNNs [97] (see Sec. 2.3), generative adversarial networks [257] and variational autoencoders [112] (see Sec. 2.5).
3.6.2 Hybrid Training
The ground state search described in section 3.1 typically starts from an NQS with randomly initialized parameters $\bm{\theta}_{0}$. Assuming convergence of the variational Monte Carlo procedure, the details of this initialization should not matter. However, if the ground state search is challenging, e.g. due to local minima, the question of convergence itself, as well as how many iterations are needed to reach the ground state can crucially depend on the choice of $\bm{\theta}_{0}$. Using existing data, e.g. from other numerical simulations or an experimental realization, the initial parameters can be chosen such that the ground state search starts from a highly promising region of the parameter space. In this case, the data is used to perform state reconstruction as described in section 3.6.1. Subsequently, the parameters of the same neural quantum state are variationally minimized to find the ground state.
In an experiment, the ground state is typically not perfectly realized. The combination of experimental data from a state close to the ground state with a subsequent numerical ground state search can yield better results than either approach exhibits on its own, as demonstrated for large, interacting two-dimensional Rydberg atom arrays using quantum Monte Carlo data in [258] and using experimental data from large two-dimensional Rydberg atom arrays [259] in [260]. In both cases, a recurrent neural network was used to represent the quantum state. In these examples, the Hamiltonian under consideration is stoquastic, and thus a wave function with real coefficients can represent its ground state. This means in particular that measurements in the computational basis are sufficient for a faithful state reconstruction.

### v3

@@ 3.3 Dynamics
Regularization schemes for $S$ have been developed, see Sec. 3.1, but these still require a very large number of samples [215] and impact the accuracy [195]. The choice of regularization method impacts the stability of the dynamics [71]. In Ref. [75], this problem is approached by disregarding the contributions to the TDVP of which there is insufficient information available due to constraints of a finite number of samples $N_{MC}$. This still requires the diagonalization of $S$, which has a high computational cost ($O(N_{\bm{\theta}}^{3})$) [215]. In Ref. [55], they adopt the minSR approach to reduce this computational cost to $O(N_{s}^{3})$, as it allows to reshape $S$ into a $N_{s}\times N_{s}$ matrix.
@@ 3.5 Open Systems
4 Learning from Data
4.1 Quantum State Tomography
@@
$\mathcal{M}(\hat{\rho})=\mathbb{E}\left[\hat{U}^{\dagger}\ket{\sigma}\bra{\sigma}\hat{U}\right]$ (58)
from measurements in the computational basis |$\ket{\sigma}$. Here, $\mathbb{E}\left[\dots\right]$ denotes the average over both the choice of unitary and the outcome distribution. The density matrix $\hat{\rho}$ can be approximated by the so-called classical shadow $\mathcal{S}$ consisting of $N_{s}$ samples, with
$\mathcal{S}=\Bigg{\{}\hat{\rho}_{i}=\mathcal{M}^{-1}\left(\hat{U}_{i}^{\dagger}\ket{\sigma}_{i}\bra{\sigma}_{i}\hat{U}_{i}\right)\quad\mathrm{with}\quad i\in 1,\dots,N_{s}\Bigg{\}}.$ (59)
@@
Also in the setting of QST, the choice of NQS depends on the state under consideration, with potential advantages e.g. of autoregressive networks and their perfect sampling [110] or of networks which can represent a high degree of entanglement [106, 266]. Typical architectures are RBMs [253, 125, 267, 37] (see Sec. 2.2), RNNs [113, 37] (see Sec. 2.5) and transformer networks [106, 266, 268] (see Sec. 2.6) and CNNs [110] (see Sec. 2.3). Furthermore, latent space representations like variational autoencoders are used [269, 270, 271]. Since these architecture has not been introduced yet, we we shortly describe autoencoders and their application to quantum state reconstruction in the following.
4.1.1 Latent Space Representations
Latent space representations consist of an encoder, the latent space or bottleneck layer, and a decoder, see Fig. 16. The encoder, typically several fully connected or convolutional layers, compresses the input into just a few nodes in the bottleneck layer. The decoder subsequently generates an output based on the information in the bottleneck layer. The network parameters are optimized such that the generated output is as close to the input as possible.
In a variational autoencoder [272], the encoder generates the values for mean and variances in the bottleneck layer, and the values used as input for the decoder are then sampled from a multivariate Gaussian.
4.2 Hybrid Training
The ground state search described in section 3.1 typically starts from an NQS with randomly initialized parameters $\bm{\theta}_{0}$. Assuming convergence of the variational Monte Carlo procedure, the details of this initialization should not matter. However, if the ground state search is challenging, e.g. due to local minima, the question of convergence itself, as well as how many iterations are needed to reach the ground state can crucially depend on the choice of $\bm{\theta}_{0}$. Using existing data, e.g. from other numerical simulations or an experimental realization, the initial parameters can be chosen such that the ground state search starts from a highly promising region of the parameter space. In this case, the data is used to perform state reconstruction as described in section 4.1. Subsequently, the parameters of the same neural quantum state are variationally minimized to find the ground state.
In an experiment, the ground state is typically not perfectly realized. The combination of experimental data from a state close to the ground state with a subsequent numerical ground state search can yield better results than either approach exhibits on its own, as demonstrated for large, interacting two-dimensional Rydberg atom arrays using quantum Monte Carlo data in [274] and using experimental snapshots from large two-dimensional Rydberg atom arrays [275] in [118]. This is done by minimizing the Kullback-Leibler (KL) divergence between the measurement distribution $q$ and the NQS amplitudes $p_{\bm{\bm{\theta}}}=|\psi_{\bm{\theta}}|^{2}$, see Eq. (57). In both cases, a recurrent neural network was used to represent the quantum state. In these examples, the Hamiltonian under consideration is stoquastic, and thus a wave function with real coefficients can represent its ground state. This means in particular that measurements in the computational basis are sufficient for a faithful state reconstruction.
@@ 4.2 Hybrid Training
The problem of the exponential cost for a reconstruction based on measurements from configurations away from the computational basis is overcome in Ref. [103] by training on observables like spin-spin correlations instead of the snapshots in the rotated basis. In order to do so, instead of the KL divergence, a variant of the mean-square error loss is used for the rotated basis. This loss function does not incorporate information on the measurement statistics, but allows to compensate for systematic errors that are e.g. present in experimental data, for example by explicitly applying spatial symmetries when calculating the experimental observables. Furthermore, for the computational basis Ref. [103] compares the KL divergence to the Wasserstein (or earth movers distance), with the advantage of the latter that it can incorporate information on the energy to the data-driven pretraining.